# Towards Vision Enhancing LLMs: Empowering Multimodal Knowledge Storage and Sharing in LLMs

### v1

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have achieved significant multimodal generation capabilities, akin to GPT-4. These models predominantly map visual information into language representation space, leveraging the vast knowledge and powerful text generation abilities of LLMs to produce multimodal instruction-following responses. We could term this method as LLMs for Vision because of its employing LLMs for visual-language understanding, yet observe that these MLLMs neglect the potential of harnessing visual knowledge to enhance overall capabilities of LLMs, which could be regraded as Vision Enhancing LLMs. In this paper, we propose an approach called MKS2, aimed at enhancing LLMs through empowering Multimodal Knowledge Storage and Sharing in LLMs. Specifically, we introduce the Modular Visual Memory, a component integrated into the internal blocks of LLMs, designed to store open-world visual information efficiently. Additionally, we present a soft Mixtures-of-Multimodal Experts architecture in LLMs to invoke multimodal knowledge collaboration during generation. Our comprehensive experiments demonstrate that MKS2 substantially augments the reasoning capabilities of LLMs in contexts necessitating physical or commonsense knowledge. It also delivers competitive results on multimodal benchmarks.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Comparisons between the proposed MKS2 and previous supervised fine-tuned (SFT) and multimodal LLMs. MKS2 focuses on improving LLMs with visual knowledge. VMN refers to the visual mapping network, transferring image encoding to the language space. MVM and MoMEs represent the proposed modular visual memory and soft mixtures-of-multimodal experts architecture in LLMs, respectively.
@@
Recent advances (Yin et al., 2023; Driess et al., 2023; Li et al., 2023c; Ye et al., 2023) on Multimodal Large Language Models (MLLMs) have opened the eyes of text-only large language models (LLM, “blind” to visual information), allowing them to understand and process multimodal information, thereby promoting the further development of LLMs-centered Artificial General Intelligence (AGI). In this line of works such as MiniGPT-4 (Zhu et al., 2023), LLaVA (Liu et al., 2023a), and BLIP-2 (Li et al., 2023a), information outside language modality is usually aligned into language space, and then the rich knowledge stored in LLM and its powerful text generation capability are used to understand various multimodal information and generate the response corresponding to human instructions. They took an significant step towards constructing a multimodal large visual-language model similar to GPT-4 (OpenAI, 2023), contributing a lot of multimodal instruction-following data (Zhang et al., 2023; Liu et al., 2023a, b) and efficient multimodal fine-tuning technical (Ye et al., 2023; Zhu et al., 2023). These approaches concentrating on multimodal information understanding could be regarded as “LLMs for vision” because of mainly utilizing LLMs for processing visual-language problems.
@@
However, current MLLMs, pretrained and supervised fine-tuned (SFT) LLMs both overlook enhancing the ability of LLMs to tap into visual knowledge. Ideally, just as the human brain retains and utilizes visual information, MLLMs or LLMs should be equipped to store external visual information. In situations that require visual common sense, even in the absence of direct visual input, LLMs should be able to access this stored visual-language knowledge for combined reasoning. This goes beyond merely processing multimodal input, as “LLMs for Vision” depicted in Figure 1. Hence, we present a term “Vision Enhancing LLMs” to describe the desired capability for LLMs. Through this enhancement, large models would store and effectively draw upon multimodal knowledge and their knowledge base and reasoning capabilities would be enhanced.
@@
To this end, We present MKS2, an innovative approach designed for empowering Multimodal Knowledge Storage and Sharing within LLM, consisting of two core stages: Visual Information Storage and Multimodal Knowledge Collaboration. In the first stage, we introduce Modular Visual Memory (MVM) in internal transformer blocks of LLMs to store visual information. Specifically, inspired by previous works (Kazemnejad et al., 2023; Wang et al., 2022b) focused on measuring parametric knowledge of pretrained language models and observing the knowledge storage role of feed-forward neural networks (FNN), we incorporate a two layers of FNN into each LLM block to build a lightweight visual memory. Subsequently, we employ a collection of image-text pairs to exclusively train and update MVM using two learning approaches: image-to-text generation and text-to-image retrieval. In both ways, soft image and token embeddings pass through the visual memory following attention calculations. These strategies empower LLMs to comprehend, translate, and store visual information in LLMs via a linguistic framework.
@@
For multimodal knowledge collaboration, we introduce a soft Mixtures-of-Multimodal Experts (MoMEs) architecture. This framework leverages specialized experts, including the Modular Visual Memory (Visual Expert) and the original MLPs (Textual Expert) in LLMs during the generation process. To efficiently achieve this, we freeze all parameters of LLMs, apply Low-Rank Adaption (LoRA (Hu et al., 2021)) to each expert module and facilitate information integration across LLM blocks through a token-level soft mixing approach. By doing so, the overall model becomes adept at accommodating both multimodal and text-modality information, enabling seamless collaboration across various input forms. During training, we collect a diverse set of instruction data, containing text-only instructions and image-text multimodal instruction-following data, to ensure the effectiveness of MoMEs in handling multimodal as well as text-only tasks.
To validate the effectiveness of our approach, we evaluate MKS2 on seven natural language processing (NLP) benchmarks and six image-text understanding datasets. Extensive experiment results indicate that MKS2 achieves superior performances on NLP tasks requiring physical or visual world knowledge, e.g, MKS2-Llama-2 significantly exceeds Llama-2-chat as shown in Figure 2. It also achieves competitive performances on image-text understanding scenarios compared to previous MLLMs.
@@
Figure 2: MKS2-Llama-2-13b achieves SOTA zero-shot performance on seven natural language reasoning tasks. It indicates achieving multimodal knowledge storage and share is effective for improving the overall capability of LLMs.
@@
  • MKS2 demonstrates superior outcomes in knowledge-intensive tasks over traditional SFT LLMs and LLMs employing Reinforcement Learning from Human Feedback (RLHF).
@@
  • Ablation studies validate the efficacy of mixtures-of-multimodal-experts that incorporates a visual knowledge expert. This architecture distinctly improves the performance of LLMs beyond the capacities of conventional supervised fine-tuned LLMs.
  • Our experiments indicate that multimodal instruction-following data further enhances LLMs’ performance in natural language reasoning tasks that require extensive commonsense.
Figure 3: The overall work flow of MKS2. It realizes visual information storage and multimodal knowledge collaboration in LLMs. In the first stage, we introduce the modular visual memory (MVM, presented in blue blocks) and train it through language-centered learning strategies. We also present a soft mixtures-of-multimodal experts (MoMEs) architecture to accomplish multimodal knowledge collaboration during generation.
@@ 2 Preliminaries
We first review the supervised fine-tuning approach and recently proposed multimodal instruction-following tuning method for LLMs.
@@ 2.1 Supervised Fine-tuning
A pure pretrained LLM is fine-tuned on high-quality labeled datasets using token-level supervision to produce a Supervised Fine-Tuned model, dubbed as SFT-LLM. Common methods are using GPT-4 automatically constructed instruction data (Wang et al., 2022c) and manually annotated high-quality data from downstream tasks (Chung et al., 2022) to fine-tune pure LLMs. To reduce training costs, recent works present some efficient instruction-tuning approaches, e.g., LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2023), etc. These SFT-LLMs are capable of generating human-like responses for various text-only instructions, having a profound impact on all walks of life.
@@ 2.2 Multimodal Instruction-Following Tuning
Compared to traditional visual-language models such as Oscar (Li et al., 2020), Flamingo (Alayrac et al., 2022), OFA (Wang et al., 2022a), etc, the multimodal instruction-following tuning approach explored extending the text-only instruction tuning in LLMs to multi-modality. These MLLMs applying LLMs as the multimodal information processor achieve impressive zero-shot performances on unseen tasks. Generally, as the traditional approach depicted in Figure 1, a frozen visual encoder (e.g., visual encoder of CLIP) is used to obtain the sequence representation of an image and a visual mapping network (VMN, a linear projection layer or Q-former from BLIP-2) projects the image encoding into soft image embeddings into the language space of LLMs. Then, we can utilize an efficient fine-tuning technical to allow LLMs to process multimodal information, thereby turning LLMs into MLLMs.
Formally, a multimodal image-text instruction sample could be expressed in the following triplet form, i.e., $(I,T,R)$, where $I,T,R$ represent the input image, text description (about human demands or image-related premises), and ground-truth response, respectively. During training, the constructed MLLMs is forced to predict the next token of response via the autoregressive objective, which could be presented as:
@@
In conclusion, we find that these two approaches ignore introducing visual knowledge to improve overall capabilities of LLMs for processing text-only tasks.
@@ 3.1 Visual Information Storage
Modular Visual Memory (MVM). This module is two layers of feed-forward neural networks (FFN) and injected into each transformer block of LLMs. As the top part shown in Figure 3, the input image $I$ is first projected into soft image embedding $\mathbf{h}_{I}$ via the pretrained visual encoder, Q-former from BLIP-2, and a learnable linear layer. Take the first block as an example; the calculation process can be presented as follows:
@@
Language-Centered Learning Strategies. As we consider LLMs as analogs to the human brain, we have embarked on a groundbreaking endeavor to create the visual storage memory in LLMs. Our ultimate goal is to empower LLMs with the capability to comprehend a given image and conjure related visual scenarios based on textual input, akin to human cognition. To this end, we adopt two learning objects to train MVM with a large mount of image-text pairs. As shown in Figure 3, we allow LLMs to generate the language description of an image, which resembles understanding and translating an image like brain. Additionally, given a sentence with some visual objects, LLM should attach to the sentence-related image, which resembles imagination. Suppose that the short description (caption) of an input image $I$ is $D$, the description generation loss is
@@
where $N$ is the number of image-text pairs in a batch and $l_{c}$ refers to the cross-entropy loss.
While retrieving the related image, we use the output hidden state $h_{e}$ of the end token $/s\right>$ of input caption to match the image embedding. Concretely, we employ a learnable linear layer to project it into the same dimension with image global encoding obtained by the visual encoder. Then we calculate the cosine similarity between them and minimize the InfoNCE loss for text-to-image (t2i) retrieval over a batch of $N$ samples. The negatives are other irrelevant images in a batch. Hence, the total language-centered learning loss is
@@
$\mathcal{L}_{t2i}=-\frac{1}{N}\sum_{i=1}^{N}\left(\log\frac{\exp\left(\operatorname{sim}\left(D_{i},IMG_{i}\right)/\tau\right)}{\sum_{j=1}^{N}\exp\left(\operatorname{sim}\left(D_{i},IMG_{j}\right)/\tau\right)}\right),$

### v2

# Vision Enhancing LLMs: Empowering Multimodal Knowledge Storage and Sharing in LLMs

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have achieved significant multimodal generation capabilities, akin to GPT-4. These models predominantly map visual information into language representation space, leveraging the vast knowledge and powerful text generation abilities of LLMs to produce multimodal instruction-following responses. We could term this method as LLMs for Vision because of its employing LLMs for visual understanding and reasoning, yet observe that these MLLMs neglect the potential of harnessing visual knowledge to enhance the overall capabilities of LLMs, which could be regarded as Vision Enhancing LLMs. In this paper, we propose an approach called MKS2, aimed at enhancing LLMs through empowering Multimodal Knowledge Storage and Sharing in LLMs. Specifically, we introduce Modular Visual Memory (MVM), a component integrated into the internal blocks of LLMs, designed to store open-world visual information efficiently. Additionally, we present a soft Mixture of Multimodal Experts (MoMEs) architecture in LLMs to invoke multimodal knowledge collaboration during text generation. Our comprehensive experiments demonstrate that MKS2 substantially augments the reasoning capabilities of LLMs in contexts necessitating physical or commonsense knowledge. It also delivers competitive results on image-text understanding multimodal benchmarks. The codes will be available at: https://github.com/HITsz-TMG/MKS2-Multimodal-Knowledge-Storage-and-Sharing
@@ 1 Introduction
Recent advances (Yin et al., 2023; Driess et al., 2023; Li et al., 2024a; Ye et al., 2023) on multimodal large language models (MLLMs) have opened the eyes of text-only large language models (LLM, “blind” to visual information), allowing them to understand and process multimodal information, thereby promoting the further development of LLMs-centered Artificial General Intelligence (AGI). In this line of works, such as MiniGPT-4 (Zhu et al., 2024), LLaVA (Liu et al., 2023a), and BLIP-2 (Li et al., 2023b), information outside language modality is usually aligned into the language space. Then, the rich knowledge stored in LLM and its powerful text generation capability are used to understand various multimodal information and generate the response corresponding to human instructions. They took a significant step towards constructing a multimodal large visual-language model similar to GPT-4 (OpenAI, 2023), contributing a lot of multimodal instruction-following data (Zhang et al., 2023b; Liu et al., 2023a, 2024c) and efficient multimodal fine-tuning technical (Ye et al., 2023; Zhu et al., 2024). These approaches, concentrating on multimodal information understanding, could be regarded as “LLMs for vision”, primarily because they mainly utilize LLMs for processing visual understanding problems.
Figure 1: Comparisons between the proposed MKS2 and previous supervised fine-tuned (SFT) and multimodal LLMs. MKS2 focuses on improving LLMs with visual knowledge. VMN refers to the visual mapping network, transferring image encoding to the language space. MVM and MoMEs represent the proposed modular visual memory and the architecture of a soft mixture of multimodal experts in LLMs, respectively.
@@
However, whether current MLLMs or pretrained and Supervised Fine-Tuned (SFT) LLMs overlook enhancing the ability of LLMs to tap into visual knowledge. Most existing multimodal models are mainly designed for vision-language understanding, focusing on interpreting or reasoning about a given image in response to a text query. They cannot acquire new, generalizable knowledge from visual data to answer future text-only questions. Similarly, text-only LLMs only learn, store, and apply knowledge from textual data. Yet, some visual knowledge cannot be fully captured by language alone, and learning from visual experience may enhance both language and object understanding in LLMs. Ideally, just as the human brain retains and utilizes visual information, MLLMs or LLMs should be equipped to store external visual information. In situations that require visual common sense, even in the absence of direct visual input, LLMs should be able to access this stored visual-language knowledge for combined reasoning. This goes beyond merely processing multimodal input, as “LLMs for Vision” depicted in the top part of Figure 1. Hence, we first present a term “Vision Enhancing LLMs" to describe the desired capability for LLMs. Through this enhancement, large models would store and effectively draw upon multimodal knowledge, and their knowledge base and reasoning capabilities would be enhanced.
@@
To this end, we present MKS2, an innovative approach designed for empowering Multimodal Knowledge Storage and Sharing within LLM, consisting of two core stages: Visual Information Storage and Multimodal Knowledge Collaboration. In the first stage, we introduce Modular Visual Memory (MVM) in internal transformer blocks of LLMs to store visual information. Specifically, inspired by previous works (Kazemnejad et al., 2023; Wang et al., 2022c) focused on measuring parametric knowledge of pretrained language models and observing the knowledge storage role of feed-forward neural networks (FNN), we incorporate two layers of FNN into each LLM block to build a lightweight visual memory. Subsequently, we employ a collection of image-text pairs to exclusively train and update MVM using two learning approaches: image-to-text generation and text-to-image retrieval or generation. The other parameters of pretrained LLMs will be frozen during training. In both ways, soft image and token embeddings pass through the visual memory following attention calculations. These learning strategies mainly empower LLMs to comprehend, translate, and store visual information via visual memory in LLMs and with a vision-centric learning approach.
@@
For multimodal knowledge collaboration, we introduce a soft Mixture of Multimodal Experts (MoMEs) architecture to recall stored visual and language knowledge. This framework leverages specialized experts, including the pretrained Modular Visual Memory (Visual Expert) and the originally trained MLPs (Textual Expert) in LLMs during the generation process. To efficiently achieve this, we freeze all parameters of LLMs, apply Low-Rank Adaption (LoRA (Hu et al., 2021)) to each expert module, and facilitate information integration across LLM blocks through a token-level soft mixing approach. By doing so, the overall model becomes adept at accommodating both multimodal and text-modality information, enabling seamless collaboration across various input forms. During training, we collect a diverse set of instruction data, containing text-only instructions and image-text multimodal instruction-following data, to ensure the effectiveness of MoMEs in handling multimodal as well as text-only tasks.
@@
Figure 2: MKS2-Llama-2-13b achieves SOTA zero-shot performance on seven natural language reasoning tasks. It indicates that achieving multimodal knowledge storage and sharing is effective for improving LLMs.
A fundamental distinction sets MKS2 apart from mainstream MLLMs, which we characterize as a shift from “LLMs for Vision" to a new paradigm we term “Vision-Enhanced LLMs." This distinction is not merely architectural but is rooted in a divergent primary objective. Traditional MLLMs (e.g., BLIP-2, LLaVA) aim to equip LLMs with visual understanding, typically evaluated on tasks like VQA. In this framework, the image is the subject of the query, and the LLM acts as a visual interpreter. While this approach has proven highly effective for multimodal benchmarks, it often leads to a well-documented trade-off: a degradation of the model’s pure language capabilities. This occurs because the intensive process of vision-language instruction tuning can distort the model’s original linguistic representations, as optimization is squarely focused on cross-modal alignment. In contrast, MKS2 is not merely to process images, but to use visual information as a source to enhance and enrich the LLM’s own internal knowledge base and linguistic capabilities. Here, the image is not the end goal but a source of knowledge from which the model learns. The objective is to absorb information from the visual world and integrate it into the LLM’s parametric memory in a language-compatible format. This allows the model to solve knowledge-intensive text-based questions that it could not answer before seeing the relevant visual data, much like a human “learning with their eyes open". While “LLMs for Vision" teach a model to see and describe, our “Vision-Enhanced LLM" paradigm teaches it to see and know.
To validate the effectiveness of our approach, we evaluate MKS2 on seven natural language processing (NLP) benchmarks and six image-text understanding datasets. Extensive experiment results indicate that MKS2 achieves superior performances on NLP tasks requiring physical or visual world knowledge, e.g., MKS2-Llama-2 significantly exceeds Llama-2-chat. It also achieves competitive performances on image-text understanding scenarios compared to MLLMs using the same supervised fine-tuning data.
@@
  • MKS2 demonstrates superior outcomes in knowledge-intensive tasks over traditional SFT LLMs and LLMs employing Reinforcement Learning from Human Feedback (RLHF), as the results shown in Figure 2.
@@
  • Ablation studies validate the efficacy of a mixture of multimodal experts that incorporate visual and language knowledge experts. This architecture distinctly improves the performance of LLMs with a few training parameters (<0.25<0.25% of LLM parameters).
  • Our analysis also indicates that the multimodal instruction-following data could enhance LLMs’ performance in commonsense reasoning tasks and text-only question-answering data can improve the open-ended visual question-answering tasks.
2 Related Work
2.1 Large Language Model
A large language model (LLM) serves as a sophisticated computational framework specifically crafted to generate human-like language and execute a wide array of natural language processing tasks. These models (Brown et al., 2020; Ouyang et al., 2022a), such as GPT-4 (OpenAI, 2023) and LLaMA (Gao et al., 2023), acquire their extensive capabilities by meticulously analyzing vast text datasets through advanced training methods, including self-supervised and semi-supervised learning techniques (Zhai et al., 2019). The pinnacle of LLM technology is represented by models employing a decoder-only transformer architecture (Vaswani, 2017), which optimizes both text processing and generation efficiency. Additionally, modern LLMs can be tailored for specific use cases or further enhanced through prompt engineering (Liu et al., 2023b), allowing users to achieve desired outputs more effectively. The training target (Gloeckle et al., 2024) for large language models is predicting the next token that follows a given sequence of tokens. By iteratively making predictions and adjusting its internal parameters based on the differences between its predicted tokens and the actual next tokens in the training data, the model learns to capture complex patterns and structures inherent in language. While the primary training corpus (Fang et al., 2024; Muennighoff et al., 2023) for these models predominantly comprises textual data sourced from the Internet, much of this text has already been utilized in the pertaining stage. However, it is important to recognize that images, videos, and other visual mediums (Zeng et al., 2023; Pu et al., 2024; Song et al., 2024) possess a wealth of common-sense knowledge that remains largely untapped. This visual information presents an opportunity to significantly augment the performance and capabilities of large language models, especially for expanding the knowledge storage of LLM. In this paper, we focus on exploring how to leverage visual knowledge derived from images to enhance the overall performance of LLMs, ultimately aiming to bridge the gap between textual and visual understanding in artificial intelligence.
2.2 Visual knowledge enhanced methods
The integration of explicit visual information has been a significant area of research aimed at enhancing the imaginative representation capabilities of language models, ultimately fostering their diverse generative potential. A notable contribution in this field is by (Jin et al., 2022; Xuan et al., 2024), who investigates how visual knowledge can be harnessed within natural language processing tasks. They develop various cross-model enhancement techniques designed to bolster the representation abilities of pretrained language models, paving the way for richer and more contextually aware text generation. Several other studies (Shi et al., 2019; Lu et al., 2022; Li et al., 2023e; Xie et al., 2024) have explored the approach of retrieving relevant images (Shi et al., 2024) from an image corpus based on textual input. The retrieved images will be fed into the language model as supplementary information. This method effectively utilizes visual knowledge to enhance performance on downstream tasks, including text completion (Zellers et al., 2019), story generation (Fan et al., 2018), and concept-to-text transformations (Barzilay & Lapata, 2005). Recently, researchers (Long et al., 2021; Yang et al., 2021; Zhu et al., 2022) have introduced innovative techniques that utilize advanced text-to-image technologies to create imaginative representations of language. These representations are then integrated into language models through prefix-tuning (Li & Liang, 2021) methodologies. Historically, approaches for enhancing language models with visual information have primarily focused on smaller models, with many methods relying on retrieval-augmented pipelines to incorporate visual knowledge. In this paper, we aim to investigate the concept of visual information storage within large language models and demonstrate how these models can be enhanced with visual knowledge without the need for explicit image inputs.
2.3 LLMs for vision
Recent works (Zhang et al., 2023b; Zhu et al., 2024; Li et al., 2023b) towards multimodal LLMs focus on utilizing the extensive knowledge and language generation capabilities of LLMs to solve multimodal tasks, especially for visual understanding and reasoning. The construction of these multimodal LLMs mainly contains the following steps. Firstly, these works usually map the visual information obtained by a pretrained visual encoder into the representation space of LLMs, through a learnable linear projection layer (Merullo et al., 2023), MLP, or Q-Former (Li et al., 2023b). This stage is usually called visual-language feature alignment and only a few data may be needed to do a good job (Liu et al., 2023a). Secondly, the initial MLLMs will be tuned via multimodal instruction-following data (Ye et al., 2023; Li et al., 2024a; Bai et al., 2023). At this stage, LLMs and the projection layer are often tuned together only with multimodal instruction data. The commonly used LLM is the SFT-LLM tuned with the lightweight LoRA (Hu et al., 2021).
In parallel, advances in instruction-tuning text-only LLMs have demonstrated remarkable performance in NLP tasks and human-machine interactions, as seen in models like Flan-T5, Bloomz (Muennighoff et al., 2022), and ChatGPT (Ouyang et al., 2022b; Muennighoff et al., 2022; Chung et al., 2022). Recently, some researchers have explored using multimodal instruction data (Liu et al., 2023a; Zhang et al., 2023a; Li et al., 2025) to fine-tune pre-trained LLMs to improve their multimodal human-machine interaction capability. (Liu et al., 2023a) employs the GPT-4 to produce the multimodal instruction data and fine-tune the language model LLaMA (Touvron et al., 2023) on the synthetic multimodal instruction-following dataset. (Zhu et al., 2024), (Li et al., 2024b), (Cha et al., 2024), and (Wang et al., 2023) also introduce a well-aligned multimodal instruction-following dataset to fine-tune a robust instruction-tuned language model (Vicuna). Recent advanced MLLMs usually achieve superior performance on open-domain multimodal question-answering tasks by integrating more powerful language models and larger, higher-quality instruction datasets, as demonstrated by models like LLaMA-VID (Li et al., 2023c), Qwen-VL (Bai et al., 2023), MiniGPT-5 (Zheng et al., 2023), InternLM-XComposer2 (Dong et al., 2024) and MobileVLM V2 (Chu et al., 2024). Moreover, some researchers are focusing on building MLLMs capable of processing high-resolution images while ensuring high inference efficiency and safety, as explored in studies such as Satety Fine-tuning (Zong et al., 2024), MoE-LLaVA (Lin et al., 2024), and LLaVA-NeXT (Liu et al., 2024b). In conclusion, these works emphasize leveraging the textual knowledge of large language models to enhance visual understanding and reasoning, contributing to the development of robust MLLMs.
@@ 3.1 Supervised Fine-tuning
A pure pretrained large language model is often fine-tuned on high-quality labeled datasets using token-level supervision to produce a Supervised Fine-Tuned model, dubbed SFT-LLM. Common methods are using GPT-4 automatically constructed instruction data (Wang et al., 2022d) and manually annotated high-quality data from downstream tasks (Chung et al., 2022) to fine-tune pure LLMs. Recent works present efficient instruction-tuning approaches to reduce training costs, e.g., LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2024), and LoftQ (Li et al., 2023d). These SFT-LLMs are capable of generating human-like responses for various text-only instructions, having a profound impact on all walks of life.
Figure 3: The overall workflow of MKS2. It realizes visual information storage and multimodal knowledge collaboration in LLMs. In the first stage, we introduce the modular visual memory (MVM) and train it through language-centric learning strategies on large-scale image-text pairs. We also present a soft mixture-of-multimodal experts (MoMEs) architecture to accomplish multimodal knowledge collaboration during text generation.
@@ 3.2 Multimodal Instruction-Following Tuning
Compared to traditional visual-language models such as Oscar (Li et al., 2020), Flamingo (Alayrac et al., 2022), OFA (Wang et al., 2022b), etc, the multimodal instruction-following tuning approach explored extending the text-only instruction tuning in LLMs to multi-modality. These MLLMs applying LLMs as the multimodal information processor achieve impressive zero-shot performances on unseen tasks. Generally, as the traditional approach depicted in Figure 1, a frozen visual encoder (e.g., the visual encoder of CLIP) is used to obtain the sequence representation of an image and a visual mapping network (VMN, a linear projection layer or Q-former from BLIP-2) projects the image encoding into soft image embeddings into the language space of LLMs. Then, we can utilize an efficient fine-tuning technique to allow LLMs to process multimodal information, thereby turning LLMs into MLLMs.
Formally, a multimodal image-text instruction sample could be expressed in the following triplet form, i.e., $(I,T,R)$, where $I,T,R$ represent the input image, text description (about human demands or image-related premises), and ground-truth response, respectively. During training, the constructed MLLMs are forced to predict the next token of response via the autoregressive objective, which could be presented as:
@@
In conclusion, we find that these two approaches overlook introducing visual knowledge to improve the overall capabilities of LLMs for processing text-only tasks.
@@ 4.1 Visual Information Storage
Modular Visual Memory (MVM). This module is a two layer of feed-forward neural networks (FFN) and is injected into each transformer block of LLMs. As the top part shown in Figure 3, the input image $I$ is first projected into soft image embedding $\mathbf{h}_{I}$ via the pretrained visual encoder (CLIP (Hafner et al., 2021)), Q-former from BLIP-2 or a simple MLP like the LLaVA (Liu et al., 2024b), and a learnable linear layer. Take the first block as an example; the calculation process can be presented as follows:
@@
Language-Centered Learning Strategies. As we consider LLMs as analogues to the human brain, we have embarked on a groundbreaking endeavour to create visual storage memory in LLMs. Our ultimate goal is to empower LLMs with the capability to comprehend a given image and conjure related visual scenarios based on textual input, akin to human cognition. To this end, we adopt two learning objects to train MVM with a large amount of image-text pairs. As the top part is shown in Figure 3, we allow LLMs to generate the language description of an image, which resembles understanding and translating an image like a brain. Additionally, given a sentence with some visual objects, LLM should attach to the sentence-related image, which resembles the imagination ability of a human. Suppose that the short description (caption) of an input image $I$ is $D$, the description generation loss is
@@
where $N$ is the number of image-text pairs in a batch and $l_{c}$ refers to the cross-entropy loss. At this stage, the visual token of the image and the text token of the caption are all passed through the newly added Visual MLP in each layer of LLM.
While retrieving the related image from sentences, we use the output hidden state $h_{e}$ of the end token $\left</s\right$> of the input caption to match the image embedding. Concretely, we employ a learnable linear layer to project it into the same dimension with image global encoding obtained by the visual encoder. Then we calculate the cosine similarity between them and minimize the InfoNCE loss for text-to-image (t2i) retrieval over a batch of $N$ samples. The negatives are other irrelevant images in a batch. Hence, the total language-centered learning loss is
@@
$\mathcal{L}_{t2i}=-\frac{1}{N}\sum_{i=1}^{N}\left(\log\frac{\exp\left(\operatorname{sim}\left(D_{i},\textit{IMG}_{i}\right)/\tau\right)}{\sum_{j=1}^{N}\exp\left(\operatorname{sim}\left(D_{i},\textit{IMG}_{j}\right)/\tau\right)}\right),$
@@
Figure 4: The detailed calculation process of the proposed soft mixture-of-multimodal experts (MoMEs) architecture. It aims to realize multimodal knowledge collaboration during text generation.